# biblio.bib

### v1

@@ Abstract
We describe a class of spin chains with new physical and computational properties. On the physical side, the spin chains give examples of symmetry-protected topological phases that are defined by non-onsite symmetries, i.e. symmetries that are not a tensor product of single-site operators. These phases can be detected by string-order parameters, but notably do not exhibit entanglement spectrum degeneracy. On the computational side, the spin chains represent a new class of states that can be used to deterministically teleport information across long distances, with the novel property that the necessary classical side processing is a non-linear function of the measurement outcomes. We also give examples of states that can serve as universal resources for measurement-based quantum computation, providing the first examples of such resources without entanglement spectrum degeneracy. The key tool in our analysis is a new kind of tensor network representation which we call split-index matrix product states (SIMPS). We develop the basic formalism of SIMPS, compare them to matrix product states, show how they are better equipped to describe certain kinds of non-onsite symmetries and quantum teleportation, and discuss how they are also well-suited to describing constrained spin chains.
@@ 1 Introduction
Quantum spin chains are an endless well of surprising physics. Models such as the Affleck-Kennedy-Lieb-Tasaki (AKLT) chain [Affleck1989], which initially appear featureless as they break no symmetries, turn out to have rich physics including degeneracy in the boundary and entanglement spectrum, hidden symmetry breaking, and non-local string order [Kennedy1992, denNijs1989, Pollmann2010, Pollmann2012a]. These properties turn out to be a consequence of the presence of symmetry-protected topological (SPT) order [Pollmann2010, Chen2011, Schuch2011, Pollmann2012]. Remarkably, these same spin chains are also useful for quantum information processing. For example, they can be used to teleport information across large distances [Popp2005, Else2012, Wahl2012], generate cat-like entanglement [Tantivasadakarn2021, Williamson2021], and even perform universal measurement-based quantum computation (MBQC) [Raussendorf2003, Miyake2010, Stephen2017].
@@
The computational and physical properties of quantum spin chains are intimately related. It has been proven that SPT order - or alternatively string order - implies quantum computational utility of the ground states of spin chains [Doherty2009, Else2012, Miller2015, Marvian2017, Stephen2017, Raussendorf2023]. Conversely, the ability to perform teleportation has been used to witness phase transitions [Verstraete2004, Verstraete2004a, Popp2005, Skrovseth2009] and to probe for SPT order in experiment [Azses2020, Smith2023]. These relations can be most clearly seen using the matrix product state (MPS) formalism [Perez-Garcia2006, Cirac2021]. Therein, the many-body wavefunction is described in terms of many local tensors whose symmetries underlie both the computational and physical properties of the MPS [Else2012, Pollmann2012]. More generally, the MPS formalism has established itself as the ultimate tool for studying 1D spin chains both analytically and numerically [Verstraete2008review, SCHOLLWOCK2011, Orus2014, Bridgeman2017, Cirac2021]. Therefore, any extension of MPS can potentially have impact on a wide range of fields and applications.
@@
In this paper, we construct a class of quantum spin chains that have novel physical and computational properties. The key concept that we introduce is a new kind of tensor network representation we call split-index matrix product states (SIMPS) (4). While equivalent to MPS in terms of which states can be efficiently captured, SIMPS offer an alternative description that we argue is essential to understanding the properties of the constructed models. For example, a key physical property of the models is that they possess symmetries which are non-onsite, i.e. not a tensor product of single-site unitaries. Compared to MPS, we show that SIMPS provide a more elegant way to identify and analyze such symmetries. Similarly, SIMPS offer a clearer picture of the computational capabilities of certain models, including ones that have previously appeared in the literature. We also show that SIMPS are well-suited to capturing constrained spin chains, including models exhibiting quantum many-body scars and Hilbert space fragmentation [Moudgalya2022].
@@
Overall, we show that the SIMPS formalism is an extension of the MPS formalism that better captures certain physical and computational properties of spin chains, clarifies some existing results in the literature, and enables the construction of new models. We summarize the key results of this paper in the next section.
@@ 1.1 Summary of results
  • In Sec. 2, we introduce the SIMPS formalism, determine when an MPS is suited to a SIMPS representations, show how to optimally convert between MPS and SIMPS representations, and prove a number of general results including a fundamental theorem of SIMPS.
  • In Sec. 3, we use SIMPS to construct a large class of spin chains that have SPT order under non-onsite symmetries. We characterize these phases in terms of string order, response to flux insertion, (a lack of) entanglement spectrum degeneracy, and by revealing the “universal fingerprints” in the entanglement structure.
@@ 2 The SIMPS representation
$|\bar{\psi}\rangle=\sum_{\begin{subarray}{c}i_{1},\dots,i_{N}\\ a,b\end{subarray}}\langle a|A^{i_{1}}\cdots A^{i_{N}}|b\rangle|a,i_{1}\cdots i_{N},b\rangle,$ (3)
@@
We remark that the SIMPS construction is very similar to double-line tensor networks [Shukla2018, Jiang2017, Williamson2016mpo], which share the feature that a given physical degree of freedom is used in more than one tensor in the network. Here, we examine this kind of representation from a more fundamental standpoint.
@@ 2.3 Examples of SIMPS
Example 3: Non-onsite symmetry breaking. For a more interesting example of a state with long-range order, consider the state,
$|\psi\rangle=\frac{1}{\sqrt{2}}(|++\cdots+\rangle+|C\rangle)\ ,$ (18)
where $|C\rangle$ is the cluster state (11). Observe that this state is invariant under the operator $\prod_{i=1}^{N}CZ_{i,i+1}$, which exchanges the two states in the superposition. Therefore, this state is associated with spontaneous breaking of this non-onsite symmetry. An MPS representation for the state can be obtained as a direct sum of the MPS representations of |$+$|+$\cdots$+$\rangle$| + + ⋯ + ⟩ and $|C\rangle$,
$A^{0}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&0\end{pmatrix},\quad A^{1}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&1&-1\end{pmatrix}.$ (19)
These matrices each have rank two. The corresponding SIMPS tensor then as $\chi=2$,
$\begin{array}[]{lll}B^{00}=\mathbb{1}&B^{01}=\mathbb{1}\\[6.0pt] B^{10}=\mathbb{1}&B^{11}=Z\ .\end{array}$ (20)
@@ 3 Non-onsite symmetries and SPT phases in SIMPS
In this section, we demonstrate on context where SIMPS are advantageous over MPS: capturing non-onsite symmetries. First, we review how onsite symmetries are capture in MPS. Then, we show how similar ideas can be used to analyze certain non-onsite symmetries in SIMPS. Using this, we define a class of SIMPS with non-onsite symmetries and demonstrate that they possess non-trivial SPT order. We then characterize some properties of SPT phases with non-onsite symmetries.
@@ 3.1 SPT order in MPS
If there is a set of symmetries $u_{g}$ that form a representation of a group $G$ such that $u_{g}u_{h}=u_{gh}$, then the corresponding operators $V_{g}$ that appear in Eq. (21) form a projective representation, $V_{g}V_{h}=\omega(g,h)V_{gh}$ for some phase factors $\omega(g,h)$. If $G$ is finite abelian, which we always assume from now on, we can instead characterize the representation by the relation $V_{g}V_{h}=\Omega(g,h)V_{h}V_{g}$ where $\Omega(g,h)=\omega(g,h)\omega(h,g)^{-1}$. The phases $\Omega(g,h)$ uniquely label the SPT phase to which the MPS belongs. If these phases are not all equal to one, then the MPS belongs to a non-trivial SPT phase protected by $G$ [Pollmann2010, Chen2011, Schuch2011].
If Eq. (21) is satisfied, it is straightforward to show that the MPS on OBC (3) is invariant under the operators $V_{g}\otimes u_{g}^{\otimes N}\otimes V_{g}^{*}$ where $V_{g}$ and $V_{g}^{*}$ act on the boundary spins where * denotes complex conjugation. The fact that the boundary spins in non-trivial SPT phases carry a projective representation implies there is a symmetry protected degeneracy at the physical boundary of the chain and in the entanglement spectrum [Pollmann2010].
@@ 3.2 Non-onsite symmetries in SIMPS
where we have defined $\tilde{B}^{ij}=e^{i\varphi_{ij}}B^{ij}$. This equation is non-trivial: It shows that the non-onsite symmetry $\mathcal{U}$ acts on a SIMPS by transforming each tensor without changing the bond dimension. Because of this, if $\mathcal{U}|\psi\rangle=|\psi\rangle$ for all system sizes, then Theorem 1 says that there must exist invertible matrices $V_{i}$ such that,
$e^{i\varphi_{ij}}B^{ij}=V_{i}B^{ij}V_{j}^{-1}$ (26)
Note that this was possible because we assumed that $\mathcal{U}$ is diagonal. More general symmetries can also correspond to SIMPS gauge transformations as long as the off-diagonal parts are a product of onsite permutation operators (e.g. $\prod_{i}X_{i}$). Therefore, we can understand certain non-onsite symmetries of SIMPS in terms of simple gauge transformations.
@@
The non-onsite symmetries can be identified immediately once the SIMPS representation of the states is written down. Conversely, it is not at all obvious to see these symmetries from the MPS representation in Eq. (15). This is another advantageous property of SIMPS when it comes to analyzing non-onsite symmetries.
The relations in Eq. (29) suggest that the SIMPS has non-trivial SPT order under the $Z_{2}\times Z_{2}$ symmetry group. While the symmetry operators commute, their representations in the virtual space of the SIMPS anticommute, and this was used to demonstrate non-trivial SPT order for the cluster state in Eq. (24). Demonstrating and characterizing this SPT order is the goal of the next section.
@@ 3.3 SPT order with non-onsite symmetries
In this section, we construct a large family of SIMPS that serve as fixed-point states for SPT phases with non-onsite symmetries. We then characterize these SIMPS in terms of string order parameters, response to flux insertion, entanglement spectrum signatures, and by characterizing the universal patterns of entanglement which we call “universal fingerprints”.
@@ 3.3.4 Universal fingerprints
3.4 Anomalous symmetries
Non-onsite symmetries are often associated with anomalies. In the present context, an anomalous symmetry can be described as a symmetry which does not permit a short-range entangled symmetric state [Chen2011a]. Given a group of non-onsite symmetries, one can calculate an invariant called a 3-cocycle [Chen2011a, Else2014]. If this 3-cocycle is non-trivial, then it can be shown that no injective MPS can be symmetric under this group [Chen2011a]. The alternatives are (a) the MPS is non-injective, meaning it has long-range order i.e. symmetry breaking, or (b) the state is not representable as an MPS with finite bond dimension, such as critical states of gapless Hamiltonians.
[…]
@@ 6.2 Quantum wire and SPT order
We are purposely vague in our meaning of “efficient” as different definitions can lead to different results. For example, under the assumption that all classical side processing of measurement outcomes is linear and byproduct operators are Pauli operators, Ref. [Hong2023] showed that teleportation implies SPT order with onsite symmetries. In contrast, the results herein violate the first assumption, since we needed non-linear classical processing to use our models for teleportation. Nevertheless, our models still have SPT order, albeit only under a generalized notion of symmetry (non-onsite symmetry). Our results therefore support Conjecture 1, while also showing how the notion of string order can and should be generalized. We expect that relaxing these assumptions further may lead to even more general notions of symmetry and SPT order, which should have interesting physical and computational ramifications.

### v2

@@ Abstract
We describe a class of spin chains with new physical and computational properties. On the physical side, the spin chains give examples of symmetry-protected topological phases that are defined by non-onsite symmetries, i.e. symmetries that are not a tensor product of single-site operators. These phases can be detected by string-order parameters, but notably do not exhibit entanglement spectrum degeneracy. On the computational side, the spin chains represent a new class of states that can be used to deterministically teleport information across long distances, with the novel property that the necessary classical side processing is a non-linear function of the measurement outcomes. We also give examples of states that can serve as universal resources for measurement-based quantum computation, providing the first examples of such resources without entanglement spectrum degeneracy. The key tool in our analysis is a new kind of tensor network representation which we call split-index matrix product states (SIMPS). We develop the basic formalism of SIMPS, compare them to matrix product states, show how they are better equipped to describe certain kinds of non-onsite symmetries including anomalous symmetries, and discuss how they are also well-suited to describing quantum teleportation and constrained spin chains.
@@ 1 Introduction
Quantum spin chains are an endless well of surprising physics. Models such as the Affleck-Kennedy-Lieb-Tasaki (AKLT) chain [Affleck1989], which initially appear featureless as they break no symmetries, turn out to have rich physics including degeneracy in the boundary and entanglement spectrum, hidden symmetry breaking, and non-local string order [Kennedy1992, denNijs1989, Pollmann2010, Pollmann2012a]. These properties are a consequence of the presence of symmetry-protected topological (SPT) order [Pollmann2010, Chen2011, Schuch2011, Pollmann2012]. Remarkably, these same spin chains are also useful for quantum information processing. For example, they can be used to teleport information across large distances [Popp2005, Else2012, Wahl2012], generate cat-like entanglement [Tantivasadakarn2021, Williamson2021], and even perform universal measurement-based quantum computation (MBQC) [Raussendorf2003, Miyake2010, Stephen2017].
@@
The computational and physical properties of quantum spin chains are intimately related. It has been proven that SPT order—or alternatively string order—implies quantum computational utility of the ground states of spin chains [Doherty2009, Else2012, Miller2015, Marvian2017, Stephen2017, Raussendorf2023]. Conversely, the ability to perform teleportation has been used to witness phase transitions [Verstraete2004, Verstraete2004a, Popp2005, Skrovseth2009] and to probe for SPT order in experiment [Azses2020, Smith2023]. These relations can be most clearly seen using the matrix product state (MPS) formalism [Perez-Garcia2006, Cirac2021]. Therein, the many-body wavefunction is described in terms of many local tensors whose symmetries underlie both the computational and physical properties of the MPS [Else2012, Pollmann2012]. More generally, the MPS formalism has established itself as the ultimate tool for studying 1D spin chains both analytically and numerically [Verstraete2008review, SCHOLLWOCK2011, Orus2014, Bridgeman2017, Cirac2021]. Therefore, any extension of MPS can potentially have impact on a wide range of fields and applications.
@@
In this paper, we construct a class of quantum spin chains that have novel physical and computational properties. The key concept that we introduce is a new kind of tensor network representation we call split-index matrix product states (SIMPS) (4). While equivalent to MPS in terms of which states can be efficiently captured, SIMPS offer an alternative description that we argue is essential to understanding the properties of the constructed models. For example, a key physical property of the models is that they possess symmetries which are non-onsite, i.e. not a tensor product of single-site unitaries. Compared to MPS, we show that SIMPS provide a much simpler way to identify and analyze such symmetries. Similarly, SIMPS offer a clearer picture of the computational capabilities of certain models, including ones that have previously appeared in the literature. We also show that SIMPS are well-suited to capturing constrained spin chains, including models exhibiting quantum many-body scars and Hilbert space fragmentation [Moudgalya2022].
@@
Overall, the SIMPS formalism is an extension of the MPS formalism that better captures certain physical and computational properties of spin chains, clarifies some existing results in the literature, and enables the construction of new models. We summarize the key results of this paper in the next section.
@@ 1.1 Summary of results
  • In Sec. 2, we introduce the SIMPS formalism, determine when an MPS is suited to a SIMPS representation, show how to optimally convert between MPS and SIMPS representations, and prove a number of general results including a fundamental theorem of SIMPS.
  • In Sec. 3, we use SIMPS to construct a large class of spin chains that have SPT order under non-onsite symmetries. We characterize these phases in terms of string order, response to flux insertion, (a lack of) entanglement spectrum degeneracy, and by revealing the “universal fingerprints” in the entanglement structure. We also show how SIMPS are well-equipped to handle anomalous symmetries.
@@ 2 The SIMPS representation
$|\bar{\psi}\rangle=\sum_{\begin{subarray}{c}i_{1},\dots,i_{N}\\ a,b\end{subarray}}\langle a|A^{i_{1}}A^{i_{2}}\cdots A^{i_{N}}|b\rangle|a,i_{1}\cdots i_{N},b\rangle,$ (3)
@@
We remark that the SIMPS construction is very similar to double-line tensor networks [Shukla2018, Jiang2017, Williamson2016mpo], which share the feature that a given physical degree of freedom is used in more than one tensor in the network (see also [Lami2022]). Here, we examine this kind of representation from a more fundamental standpoint.
@@ 2.3 Examples of SIMPS
Example 3: State with anomalous symmetry. For a more interesting example of a state with long-range order, consider the following state from Ref. [GarreRubio2023],
$|\psi\rangle=\frac{1}{\sqrt{2}}(|++\cdots+\rangle+\prod_{i}Z_{i}CZ_{i,i+1}|++\cdots+\rangle)\ .$ (18)
This state is invariant under the operator $U_{CZX}=\prod_{i}X_{i}\prod_{i=1}^{N}Z_{i}CZ_{i,i+1}$ which is an anomalous $Z_{2}$ symmetry [Chen2011, GarreRubio2023]. An MPS representation for the state can be obtained as a direct sum of the MPS representations of |$+$|+$\cdots$+$\rangle$| + + ⋯ + ⟩ and $\prod_{i}Z_{i}CZ_{i,i+1}|++\cdots+\rangle)$,
$A^{0}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&0\end{pmatrix},\quad A^{1}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&-1&1\end{pmatrix}.$ (19)
These matrices each have rank two. The corresponding SIMPS tensor then has $\chi=2$,
$\begin{array}[]{lll}B^{00}=\mathbb{1}&B^{01}=\mathbb{1}\\[6.0pt] B^{10}=Z&B^{11}=\mathbb{1}\ .\end{array}$ (20)
@@ 3 Non-onsite symmetries and SPT phases in SIMPS
In this section, we demonstrate the one context where SIMPS are advantageous over MPS: capturing non-onsite symmetries. First, we review how onsite symmetries are capture in MPS. Then, we show how similar ideas can be used to analyze certain non-onsite symmetries in SIMPS. This includes a general class of anomalous group-like symmetries. We then define a class of SIMPS possessing non-trivial SPT order protected by non-onsite symmetries and characterize their physical properties.
@@ 3.1 SPT order in MPS
If there is a set of symmetries $u_{g}$ that form a representation of a group $G$ such that $u_{g}u_{h}=u_{gh}$, then the corresponding operators $V_{g}$ that appear in Eq. (21) form a projective representation, $V_{g}V_{h}=\omega(g,h)V_{gh}$ for some phase factors $\omega(g,h)$ [Schuch2011]. If $G$ is finite abelian, which we always assume from now on, we can instead characterize the representation by the relation $V_{g}V_{h}=\Omega(g,h)V_{h}V_{g}$ where $\Omega(g,h)=\omega(g,h)\omega(h,g)^{-1}$. The phases $\Omega(g,h)$ uniquely label the SPT phase to which the MPS belongs [Berkovich1998]. If these phases are not all equal to one, then the MPS belongs to a non-trivial SPT phase protected by $G$ [Pollmann2010, Chen2011, Schuch2011].
If Eq. (21) is satisfied, it is straightforward to show that the MPS on OBC (3) is invariant under the operators $V_{g}\otimes u_{g}^{\otimes N}\otimes V_{g}^{*}$ where $V_{g}$ and $V_{g}^{*}$ act on the boundary spins where * denotes complex conjugation. The fact that the boundary spins in non-trivial SPT phases carry a projective representation implies a symmetry protected degeneracy at the physical boundary of the chain and in the entanglement spectrum [Pollmann2010].
@@ 3.2 Non-onsite symmetries in SIMPS
where we have defined $\tilde{B}^{ij}=e^{i\varphi_{ij}}B^{ij}$. This equation is non-trivial: It shows that the non-onsite symmetry $\mathcal{U}$ acts on a SIMPS by transforming each tensor without changing the bond dimension. This is possible since the split-index structure effectively splits the different gates $U$ across different tensors. Because of this, if $\mathcal{U}|\psi\rangle=|\psi\rangle$ for all system sizes and the SIMPS is normal, then Theorem 1 says that there must exist invertible matrices $V_{i}$ such that,
$e^{i\varphi_{ij}}B^{ij}=V_{i}B^{ij}V_{j}^{-1}\ .$ (26)
Note that this was possible because we assumed that $\mathcal{U}$ is diagonal. More general symmetries can also correspond to SIMPS gauge transformations as long as the off-diagonal parts are a product of onsite permutation operators (e.g. $\prod_{i}X_{i}$) since such operators can also be split to act on each tensor independently. Therefore, we can understand certain non-onsite symmetries of SIMPS in terms of simple gauge transformations. We discuss this class of symmetries in more detail in Sec. 3.3.
@@
In this example, the non-onsite symmetries could be identified immediately once the SIMPS representation of the states is written down. Conversely, it is not at all obvious to see these symmetries from the MPS representation in Eq. (15). This is another advantageous property of SIMPS when it comes to analyzing non-onsite symmetries.
The relations in Eq. (29) suggest that the SIMPS has non-trivial SPT order under the $Z_{2}\times Z_{2}$ symmetry group. While the symmetry operators commute, their representations in the virtual space of the SIMPS anticommute, and this was used to demonstrate non-trivial SPT order for the cluster state in Eq. (24). Demonstrating and characterizing this SPT order is the goal of Sec. 3.4.
3.3 Anomalous symmetries
Non-onsite symmetries are often associated with anomalies, and here we show that SIMPS are well-suited to capturing anomalous symmetries. In the present context, an anomalous symmetry can be described as a symmetry which does not permit a short-range entangled symmetric state [Chen2011a]. Given a group of non-onsite symmetries, one can calculate an invariant called a 3-cocycle [Chen2011a, Else2014]. If this 3-cocycle belongs to a non-trivial equivalence class, then it can be shown that no injective MPS can be symmetric under this group [Chen2011a]. The alternatives are (a) the MPS is non-injective, meaning it has long-range order i.e. symmetry breaking, or (b) the state is not representable as an MPS with finite bond dimension, such as critical states of gapless Hamiltonians.
To demonstrate the ability of SIMPS to capture anomalous symmetries, we now write down a representative set of non-onsite symmetries for every finite group $G$ and 3-cocycle $\omega$ such that the symmetries can be captured by SIMPS gauge transformations. The representations, described in Ref. [GarreRubio2023], are defined with respect to an onsite Hilbert space $\mathbb{C}^{|G|}$ spanned by the states $|g\rangle$ for $g\in G$. Define the operators $L^{g}$ such that $L^{g}|h\rangle=|gh\rangle$ and the diagonal two-site operators $W^{g}$ such that $W^{g}|hk\rangle=\omega(g,k,k^{-1}h)|hk\rangle$. Then, the non-onsite symmetries are defined for each $g\in G$ as,
$T^{g}=\prod_{i}L^{g}_{i}\prod_{i}W^{g}_{i,i+1}\ .$ (31)
One can check that these non-onsite operators, as a representation of $G$, are characterized by the 3-cocycle $\omega$ [GarreRubio2023]. This operator consists of a product of diagonal two-site operators and single-site permutations, and therefore it can be captured via SIMPS gauge transformations. Specifically, acting with this symmetry on a SIMPS transforms the tensor in the following way,
$B^{h,k}\xrightarrow[]{T^{g}}\omega(g,k,k^{-1}h)B^{gh,gk}\ .$ (32)
If a normal SIMPS is invariant under this symmetry, the fundamental theorem (10) says that there exist invertible matrices $V^{g}_{h}$ such that,
$\omega(g,k,k^{-1}h)B^{gh,gk}=V^{g}_{h}B^{h,k}{V^{g}_{k}}^{-1}\ .$ (33)
[…]
$B^{i,j}\xrightarrow[]{U_{CZX}}(-1)^{i+ij}B^{i+1,j+1}\ .$ (34)
which is indeed equivalent to a gauge transformation for the tensor defined in Eq. (20),
$(-1)^{i+ij}B^{i+1,j+1}=(Z^{i}X)B^{i,j}(Z^{j}X)^{-1}\ .$ (35)
Therefore, we expect that SIMPS will be a useful tool for studying anomalous symmetries in future.
@@ 3.4 SPT order with non-onsite symmetries
In this section, we construct a large family of SIMPS that serve as fixed-point states for SPT phases with (non-anomalous) non-onsite symmetries. We then characterize these SIMPS in terms of string order parameters, response to flux insertion, entanglement spectrum signatures, and by characterizing the universal patterns of entanglement which we call “universal fingerprints”.
@@ 6.2 Quantum wire and SPT order
We are purposely vague in our meaning of “efficient” as different definitions can lead to different results. For example, under the assumption that all classical side processing of measurement outcomes is linear and byproduct operators are Pauli operators, Ref. [Hong2023] showed that teleportation implies SPT order with onsite symmetries. In contrast, the results herein violate the first assumption, since we needed non-linear classical processing to use our models for teleportation. Nevertheless, our models still have SPT order, albeit only under a generalized notion of symmetry (non-onsite symmetry). Our results therefore support Conjecture 1, while also showing how the notion of SPT order can and should be generalized. We expect that relaxing these assumptions further may lead to even more general notions of symmetry and SPT order, which should have interesting physical and computational ramifications.